Determine whether or not each equation is that of an ellipse, if it is not, state the kind of graph the equation has:
a. x^2+4y^2=4
b. x^2+y^2=4
c. x^2+y=4
d. x/4+y/25=1

Answer by ikleyn(52803)

.
Determine whether or not each equation is that of an ellipse, if it is not, state the kind of graph the equation has:

a. x^2+4y^2=4        ellipse

b. x^2+y^2=4         circle (a particular case of an ellipse)

c. x^2+y=4           parabola

d. x/4+y/25=1        straight line